# Saumili Janaa,1, John Kolinskib,2, Detlef Lohsea,c,3, and Vatsal Sanjaye,4

### v1

@@ 4.3 Predictive Interpolating Model for Maximum Impact Force
For intermediate conditions ($El\sim\mathcal{O}\left(1\right)$), the sphere’s deformation and the fluid’s inertia both contribute, and the peak force deviates from either pure Hertz or Wagner scaling alone. We therefore express the dimensionless peak force as a weighted transition between the two asymptotic contributions following the approach of Sanjay & Lohse (2025)5,
@@
with a smooth transition function $\text{f}\left(El\right)$ defined as a tanh function in $\log_{10}El$, centered at $El\sim\mathcal{O}\left(1\right)$. The coefficients are fixed a priori: the Hertz prefactor 5.3 follows directly from the elastic analysis above (without fitting), while the constants 3.24 (Wagner plateau) and 3.2 (the $We^{-1}$ capillary correction) are taken from the Newtonian impact model of Sanjay & Lohse (2025)5. This construction ensures a continuous interpolation between the Hertz and Wagner limits. In spirit it follows the additive scaling approach of Sanjay & Lohse (2025)5 for drop impacts, but unlike their model – which includes separate viscous regimes – here only the two primary regimes (elastic vs. inertial) are needed. The resulting formula smoothly bridges the two asymptotes and correctly reproduces the peak-force scaling in both limits (this predictive curve is plotted in fig. 4a for comparison).
@@ 5 Influence of Elastic Stress Relaxation
The results in § 3 established the two asymptotic force scalings for impacts with permanent memory ($De\to\infty$): a Wagner plateau at small elasticity number and a Hertz law at large elasticity number (figure 4). These two limits also bound impacts when the material memory is finite: relaxing the memory shifts the response continuously from solid‑like to liquid‑like, with the transition controlled by the non‑dimensional relaxation time. We quantify the memory effect with the Weissenberg number
$Wi=\frac{\lambda}{R_{0}/V_{0}}=De\sqrt{We},$ (31)
which compares the elastic relaxation time $\lambda$ to the impact time $R_{0}/V_{0}$ (see § 2.1). While $De$ compares $\lambda$ to the inertio‑capillary time, $Wi$ is the more natural process‑time measure here and increases with the degree to which elastic stresses persist during impact 46. Thus $Wi=0$ ($De=0$) recovers a Newtonian liquid, whereas $Wi\to\infty$ ($De\to\infty$) yields an elastic‑memory limit.
@@
The modeling choices in this work were made to isolate the physics of the liquid–to–elastic transition and enable direct comparison to the classical limits. The viscoelastic sphere obeys Oldroyd‑B, which neglects finite microstructure extensibility and shear‑thinning; this is acceptable for our flow history but can be systematically relaxed with constitutive models that incorporate finite extensibility and rate‑dependent viscosity 21, 32, 35, 34. The substrate is non‑contacting, so the thin gas layer is present but not explicitly resolved with lubrication and wetting dynamics; prior work shows that air cushioning, skating on a gas film, and nanoscale first contact depend sensitively on slip, compressibility and rarefaction 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57. Furthermore, the onset of contact will modify the shear stress at the interface - including effects such as adhesion - that can modify the peak force, and will alter the stresses upon rebound. Finally, a small background viscosity renders the ($De\to\infty$) limit to be of Kelvin–Voigt type rather than ideally elastic; classical analyses quantify how viscoelastic dissipation and elastic waves perturb Hertzian impact 7, 8. These assumptions do not alter the governing exponents (Wagner versus Hertz) but they may affect quantitative prefactors and very‑short‑time, near‑contact details.
Several direct extensions can sharpen and generalize these results. Experiments with soft hydrogel/elastomer beads at moderate $V_{0}$ can probe the crossover regime and test the full $F(t)$ waveform (peak magnitude, rise time, symmetry/skewness), leveraging recent studies on soft‑solid and gel impacts and elastohydrodynamic bouncing 58, 59, 60, 61, 62, 63, 9, and using established force‑measurement protocols from liquid‑drop impacts 64, 65, 66, 3, 4. Furthermore, incorporating a thin‑gas lubrication model with dynamic wetting (including slip and, if needed, rarefaction/compressibility) will resolve when and how contact initiates and how this feeds back on the very‑early‑time force 43, 47, 49, 50, 52, 51, 67, 53, 54, 57. Linking impact memory with lubricated‑impact and wetting transitions will place the near‑contact force history on firmer ground 68, 69, 70, 71, 72.
@@ Acknowledgements
Notes and references
  • Wagner 1932 H. Wagner, Über Stoß- und Gleitvorgänge an der Oberfläche von Flüssigkeiten, Z. Angew. Math. Mech., 1932, 12, 193–215.
  • Philippi et al. 2016 J. Philippi, P.-Y. Lagrée and A. Antkowiak, Drop impact on a solid surface: short-time self-similarity, J. Fluid Mech., 2016, 795, 96–135.
  • Zhang et al. 2022 B. Zhang, V. Sanjay, S. Shi, Y. Zhao, C. Lv and D. Lohse, Impact forces of water drops falling on superhydrophobic surfaces, Phys. Rev. Lett., 2022, 129, 104501.
  • Sanjay et al. 2025 V. Sanjay, B. Zhang, C. Lv and D. Lohse, The Role of Viscosity on Drop Impact Forces on Non-Wetting Surfaces, J. Fluid Mech., 2025, 1004, A6.
  • Sanjay and Lohse 2025 V. Sanjay and D. Lohse, Unifying Theory of Scaling in Drop Impact: Forces and Maximum Spreading Diameter, 2025, 134, 104003.
  • Hertz 1882 H. Hertz, Ueber die Berührung fester elastischer Körper., 1882, https://www.degruyter.com/document/doi/10.1515/crll.1882.92.156/html.
  • 7 Y.-H. Pao, Extension of the Hertz Theory of Impact to the Viscoelastic Case, 26, 1083–1088.
  • 8 S. C. Hunter, Energy Absorbed by Elastic Waves during Impact, 5, 162–171.
  • Bertin 2024 V. Bertin, Similarity solutions in elastohydrodynamic bouncing, 2024, 986, A13.
  • Ahmad et al. 2013 M. Ahmad, M. Schatz and M. Casey, Experimental investigation of droplet size influence on low pressure steam turbine blade erosion, Wear, 2013, 303, 83–86.
  • Cheng et al. 2022 X. Cheng, T.-P. Sun and L. Gordillo, Drop impact dynamics: impact force and stress distributions, Annu. Rev. Fluid Mech., 2022, 54, 57–81.
  • Lohse 2022 D. Lohse, Fundamental fluid dynamics challenges in inkjet printing, Annu. Rev. Fluid Mech., 2022, 54, 349–382.
  • Shiri and Bird 2017 S. Shiri and J. C. Bird, Heat exchange between a bouncing drop and a superhydrophobic substrate, Proc. Natl. Acad. Sci. USA, 2017, 114, 6930–6935.
  • Kim 2007 J. Kim, Spray cooling heat transfer: The state of the art, Int. J. Heat Fluid Flow, 2007, 28, 753–767.
  • Smith et al. 2018 F. R. Smith, C. Nicloux and D. Brutin, Influence of the impact energy on the pattern of blood drip stains, Phys. Rev. Fluids, 2018, 3, 013601.
  • He et al. 2021 L. He, L. Ding, B. Li, W. Mu, P. Li and F. Liu, Optimization Strategy to Inhibit Droplets Rebound on Pathogen-Modified Hydrophobic Surfaces, ACS Appl. Mater. Interfaces, 2021, 13, 38018–38028.
  • Nearing et al. 1986 M. A. Nearing, J. M. Bradford and R. D. Holtz, Measurement of force vs. time relations for waterdrop impact, Soil Sci. Soc. Am. J., 1986, 50, 1532–1536.
  • Tirupataiah and Sundararajan 1991 Y. Tirupataiah and G. Sundararajan, A dynamic indentation technique for the characterization of the high strain rate plastic flow behaviour of ductile metals and alloys, J. Mech. Phys. Solids, 1991, 39, 243–271.
  • Andreotti et al. 2013 B. Andreotti, Y. Forterre and O. Pouliquen, Granular Media: Between Fluid and Solid, Cambridge University Press, 2013.
  • Cross 1999 R. Cross, The bounce of a ball, Am. J. Phys., 1999, 67, 222–227.
  • Snoeijer et al. 2020 J. H. Snoeijer, A. Pandey, M. A. Herrada and J. Eggers, The relationship between viscoelasticity and elasticity, Proc. R. Soc. A, 2020, 476, 20200419.
  • CLASEN et al. 2006 C. CLASEN, J. EGGERS, M. A. FONTELOS, J. LI and G. H. McKINLEY, The beads-on-string structure of viscoelastic threads, J. Fluid Mech., 2006, 556, 283–308.
  • Tamim and Bostwick 2021 S. Tamim and J. Bostwick, Oscillations of a soft viscoelastic drop, npj Microgravity, 2021, 7, 42.
  • Calvert 2001 P. Calvert, Inkjet Printing for Materials and Devices, Chem. Mater., 2001, 13, 3299–3305.
  • Bergeron et al. 2000 V. B. Bergeron, D. B. Bonn, J. M. Martin and V. L. Louis, Controlling Droplet Deposition with Polymer Additives, Nature, 2000, 405, 772–775.
  • Thompson and Rothstein 2007 J. Thompson and J. Rothstein, The atomization of viscoelastic fluids in flat-fan and hollow-cone spray nozzles, J. Non-Newtonian Fluid Mech., 2007, 147, 11–22.
  • Derby 2012 B. Derby, Printing and Prototyping of Tissues and Scaffolds, Science, 2012, 338, 921–6.
  • Li and Mooney 2016 J. Li and D. Mooney, Designing hydrogels for controlled drug delivery, Nature Rev. Mater., 2016, 1, 16071.
  • Chen et al. 2010 D. Chen, Q. Wen, P. Janmey, J. Crocker and A. Yodh, Rheology of Soft Materials, Annu. Rev. Condens. Matter Phys, 2010, 1, 301–22.
  • Oldroyd and Wilson 1950 J. G. Oldroyd and A. H. Wilson, On the formulation of rheological equations of state, Proc. R. Soc. Lond. A, 1950, 200, 523–541.
  • Dixit et al. 2025 A. K. Dixit, A. Oratis, K. Zinelis, D. Lohse and V. Sanjay, Viscoelastic Worthington Jets and Droplets Produced by Bursting Bubbles, J. Fluid Mech., 2025, 1010, A2.
  • McKinley and Sridhar 2002 G. H. McKinley and T. Sridhar, Filament-stretching rheometry of complex fluids, Annu. Rev. Fluid Mech., 2002, 34, 375–415.
  • Bird et al. 1980 R. Bird, P. Dotson and N. Johnson, Polymer solution rheology based on a finitely extensible bead—spring chain model, Journal of Non-Newtonian Fluid Mechanics, 1980, 7, 213–235.
  • Tanner 2000 R. Tanner, Engineering Rheology, OUP Oxford, 2000.
  • Thien and Tanner 1977 N. P. Thien and R. I. Tanner, A new constitutive equation derived from network theory, J. Non-Newton. Fluid, 1977, 2, 353–365.
  • Popinet and collaborators 2013–2023 S. Popinet and collaborators, Basilisk C: volume of fluid method, http://basilisk.fr (Last accessed: August 23, 2023), 2013–2023.
  • Tryggvason et al. 2011 G. Tryggvason, R. Scardovelli and S. Zaleski, Direct Numerical Simulations of Gas–Liquid Multiphase Flows, Cambridge University Press, 2011.
  • Brackbill et al. 1992 J. U. Brackbill, D. B. Kothe and C. Zemach, A continuum method for modeling surface tension, J. Comput. Phys., 1992, 100, 335–354.
  • Popinet 2018 S. Popinet, Numerical models of surface tension, Annu. Rev. Fluid Mech., 2018, 50, 49–75.
  • Popinet 2009 S. Popinet, An accurate adaptive solver for surface-tension-driven interfacial flows, J. Comput. Phys., 2009, 228, 5838–5866.
  • Kolinski et al. 2014 J. M. Kolinski, L. Mahadevan and S. M. Rubinstein, Drops can bounce from perfectly hydrophilic surfaces, Europhys. Lett., 2014, 108, 24001.
  • Sanjay 2022 V. Sanjay, PhD Thesis - Research UT, graduation UT, University of Twente, Netherlands, 2022.
  • Zheng et al. 2021 S. Zheng, S. Dillavou and J. M. Kolinski, Air mediates the impact of a compliant hemisphere on a rigid smooth surface, Soft Matter, 2021, 17, 3813–3819.
  • Popinet 2015 S. Popinet, A quadtree-adaptive multigrid solver for the Serre–Green–Naghdi equations, J. Comput. Phys., 2015, 302, 336–358.
  • Negus 2022 M. Negus, PhD thesis, University of Oxford, 2022.
  • Poole 2012 R. J. Poole, The Deborah and Weissenberg numbers, Rheol. Bull., 2012, 53, 32–39.
  • Smith et al. 2003 F. T. Smith, L. Li and G. X. Wu, Air cushioning with a lubrication/inviscid balance, J. Fluid Mech., 2003, 482, 291–318.
  • Thoroddsen et al. 2003 S. T. Thoroddsen, T. G. Etoh and K. Takehara, Air entrapment under an impacting drop, J. Fluid Mech., 2003, 478, 125–134.
  • Kolinski et al. 2012 J. M. Kolinski, S. M. Rubinstein, S. Mandre, M. P. Brenner, D. A. Weitz and L. Mahadevan, Skating on a film of air: drops impacting on a surface, Phys. Rev. Lett., 2012, 108, 074503.
  • van der Veen et al. 2012 R. C. A. van der Veen, T. Tran, D. Lohse and C. Sun, Direct measurements of air layer profiles under impacting droplets using high-speed color interferometry, Phys. Rev. E, 2012, 85, 026315.
  • Driscoll and Nagel 2011 M. M. Driscoll and S. R. Nagel, Ultrafast interference imaging of air in splashing dynamics, Phys. Rev. Lett., 2011, 107, 154502.
  • Li et al. 2015 E. Q. Li, I. U. Vakarelski and S. T. Thoroddsen, Probing the nanoscale: the first contact of an impacting drop, J. Fluid Mech., 2015, 785, R2.
  • Chubynsky et al. 2020 M. V. Chubynsky, K. I. Belousov, D. A. Lockerby and J. E. Sprittles, Bouncing off the walls: the influence of gas-kinetic and van der Waals effects in drop impact, Phys. Rev. Lett., 2020, 124, 084501.
  • García-Geijo et al. 2024 P. García-Geijo, G. Riboux and J. Gordillo, The skating of drops impacting over gas or vapour layers, J. Fluid Mech., 2024, 980, A35.
  • Langley et al. 2017 K. Langley, E. Q. Li and S. T. Thoroddsen, Impact of ultra-viscous drops: air-film gliding and extreme wetting, J. Fluid Mech., 2017, 813, 647–666.
  • Zhang et al. 2021 L. Zhang, T. Soori, A. Rokoni, A. Kaminski and Y. Sun, Thin film instability driven dimple mode of air film failure during drop impact on smooth surfaces, Phys. Rev. Fluids, 2021, 6, 044002.
  • Sprittles 2024 J. E. Sprittles, Gas microfilms in droplet dynamics: When do drops bounce?, Annu. Rev. Fluid Mech., 2024, 56, 91–118.
  • Mitra et al. 2021 S. Mitra, Q. Vo and T. Tran, Bouncing-to-wetting transition of water droplets impacting soft solids, 2021.
  • Jalaal et al. 2018 M. Jalaal, C. Seyfert, B. Stoeber and N. J. Balmforth, Gel-controlled droplet spreading, J. Fluid Mech., 2018, 837, 115–128.
  • Luu and Forterre 2009 L.-H. Luu and Y. Forterre, Drop impact of yield-stress fluids, J. Fluid Mech., 2009, 632, 301–327.
  • Martouzet et al. 2021 G. Martouzet, L. Jørgensen, Y. Pelet, A.-L. Biance and C. Barentin, Dynamic arrest during the spreading of a yield stress fluid drop, Phys. Rev. Fluids, 2021, 6, 044006.
  • Chevy et al. 2012 F. Chevy, A. Chepelianskii, D. Quéré and E. Raphaël, Liquid Hertz contact: softness of weakly deformed drops on non-wetting substrates, Europhys. Lett., 2012, 100, 54002.
  • Agüero et al. 2022 E. A. Agüero, L. F. L. Alventosa, D. M. Harris and C. A. Galeano-Rios, Impact of a rigid sphere onto an elastic membrane, Proc. R. Soc. Lond. A, 2022, 478, 20220340.
  • Mitchell et al. 2019 B. R. Mitchell, J. C. Klewicki, Y. P. Korkolis and B. L. Kinsey, The transient force profile of low-speed droplet impact: measurements and model, J. Fluid Mech., 2019, 867, 300–322.
  • Li et al. 2014 J. Li, B. Zhang, P. Guo and Q. Lv, Impact force of a low speed water droplet colliding on a solid surface, J. Appl. Phys., 2014, 116, 214903.
  • Zhang et al. 2017 B. Zhang, J. Li, P. Guo and Q. Lv, Experimental studies on the effect of Reynolds and Weber numbers on the impact forces of low-speed droplets colliding with a solid surface, Exp. Fluids, 2017, 58, 1–12.
  • Mandre et al. 2009 S. Mandre, M. Mani and M. P. Brenner, Precursors to splashing of liquid droplets on a solid surface, Phys. Rev. Lett., 2009, 102, 134502.
  • Snoeijer and Andreotti 2013 J. H. Snoeijer and B. Andreotti, Moving contact lines: scales, regimes, and dynamical transitions, Annu. Rev. Fluid Mech., 2013, 45, 269–292.
  • Richard and Quéré 2000 D. Richard and D. Quéré, Bouncing water drops, Europhys. Lett., 2000, 50, 769–775.
  • Sharma and Dixit 2021 P. K. Sharma and H. N. Dixit, Regimes of wettability-dependent and wettability-independent bouncing of a drop on a solid surface, J. Fluid Mech., 2021, 908, A37.
  • Sanjay et al. 2023 V. Sanjay, P. Chantelot and D. Lohse, When does an impacting drop stop bouncing?, J. Fluid Mech., 2023, 958, A26.
  • Chantelot and Lohse 2021 P. Chantelot and D. Lohse, Drop impact on superheated surfaces: short-time dynamics and transition to contact, J. Fluid Mech., 2021, 928, A36.

### v2

@@ 4.3 Predictive Interpolating Model for Maximum Impact Force
For intermediate conditions ($El\sim\mathcal{O}\left(1\right)$), the sphere’s deformation and the fluid’s inertia both contribute, and the peak force deviates from either pure Hertz or Wagner scaling alone. We therefore express the dimensionless peak force as a weighted transition between the two asymptotic contributions following the approach of Sanjay & Lohse (2025)sanjayUnifyingTheoryScaling2025,
@@
with a smooth transition function $\text{f}\left(El\right)$ defined as a tanh function in $\log_{10}El$, centered at $El\sim\mathcal{O}\left(1\right)$. The coefficients are fixed a priori: the Hertz prefactor 5.3 follows directly from the elastic analysis above (without fitting), while the constants 3.24 (Wagner plateau) and 3.2 (the $We^{-1}$ capillary correction) are taken from the Newtonian impact model of Sanjay & Lohse (2025)sanjayUnifyingTheoryScaling2025. This construction ensures a continuous interpolation between the Hertz and Wagner limits. In spirit it follows the additive scaling approach of Sanjay & Lohse (2025)sanjayUnifyingTheoryScaling2025 for drop impacts, but unlike their model – which includes separate viscous regimes – here only the two primary regimes (elastic vs. inertial) are needed. The resulting formula smoothly bridges the two asymptotes and correctly reproduces the peak-force scaling in both limits (this predictive curve is plotted in fig. 4a for comparison).
@@ 5 Influence of Elastic Stress Relaxation
The results in § 3 established the two asymptotic force scalings for impacts with permanent memory ($De\to\infty$): a Wagner plateau at small elasticity number and a Hertz law at large elasticity number (figure 4). These two limits also bound impacts when the material memory is finite: relaxing the memory shifts the response continuously from solid‑like to liquid‑like, with the transition controlled by the non‑dimensional relaxation time. We quantify the memory effect with the Weissenberg number, $Wi$ (eq. 8), which compares the elastic relaxation time $\lambda$ to the impact time $R_{0}/V_{0}$ (see § 2.1). While $De$ compares $\lambda$ to the inertio‑capillary time, $Wi$ is the more natural process‑time measure here and increases with the degree to which elastic stresses persist during impact WiandDe_Poole. Thus $Wi=0$ ($De=0$) recovers a Newtonian liquid, whereas $Wi\to\infty$ ($De\to\infty$) yields an elastic‑memory limit.
@@
The modeling choices in this work were made to isolate the physics of the liquid–to–elastic transition and enable direct comparison to the classical limits. The viscoelastic sphere obeys Oldroyd‑B, which neglects finite microstructure extensibility and shear‑thinning; this is acceptable for our flow history but can be systematically relaxed with constitutive models that incorporate finite extensibility and rate‑dependent viscosity Snoeijer_etal_2020, Mckinley_exten_flow, Thien_1977, tanner2000. The substrate is non‑contacting, so the thin gas layer is present but not explicitly resolved with lubrication and wetting dynamics; prior work shows that air cushioning, skating on a gas film, and nanoscale first contact depend sensitively on slip, compressibility and rarefaction smith-2003-jfm, thoroddsen-2003-jfm, thoroddsen_etoh_takehara_ootsuka_hatsuki_2005, kolinski-2012-prl, vdveen-2012-pre, Driscoll2011, li-2015-jfm, chubynsky-2020-prl, garcia2024skating, langley2017impact, zhang-2021-prf, sprittles2024gas. We stress that in the present work, the substrate is assumed to remain perfectly rigid. Small elastic deformations of the substrate can, however, modify the intervening gas layer Langley2020, which in turn may alter the impact dynamics and the measured force signature of the impacting sphere.Ma Furthermore, the onset of contact will modify the shear stress at the interface - including effects such as adhesion - that can modify the peak force, and will alter the stresses upon rebound. Finally, a small background viscosity renders the ($De\to\infty$) limit to be of Kelvin–Voigt type rather than ideally elastic; classical analyses quantify how viscoelastic dissipation and elastic waves perturb Hertzian impact pao1955extension, hunterEnergyAbsorbedElastic1957. These assumptions do not alter the governing exponents (Wagner versus Hertz) but they may affect quantitative prefactors and very‑short‑time, near‑contact details.
Several direct extensions can sharpen and generalize these results. Experiments with soft hydrogel/elastomer beads at moderate $V_{0}$ can probe the crossover regime and test the full $F(t)$ waveform (peak magnitude, rise time, symmetry/skewness), leveraging recent studies on soft‑solid and gel impacts and elastohydrodynamic bouncing mitra2021bouncing, jalaal-2018-jfm, luu-2009-jfm, martouzet2021dynamic, chevy2012liquid, aguero2022impact, bertinSimilaritySolutionsElastohydrodynamic2024, and using established force‑measurement protocols from liquid‑drop impacts Mitchell2019, Li2014, Zhang2017, zhangImpactForcesWater2022, sanjayRoleViscosityDrop2025. Furthermore, incorporating a thin‑gas lubrication model with dynamic wetting (including slip and, if needed, rarefaction/compressibility) will resolve when and how contact initiates and how this feeds back on the very‑early‑time force zheng2021air, smith-2003-jfm, kolinski-2012-prl, vdveen-2012-pre, li-2015-jfm, Driscoll2011, Mandre_etalPRL_2019, chubynsky-2020-prl, garcia2024skating, sprittles2024gas. Linking impact memory with lubricated‑impact and wetting transitions will place the near‑contact force history on firmer ground snoeijer-2013-arfm, Richard2000, sharma-2021-jfm, sanjay_chantelot_lohse_2023, chantelot_lohse_2021.